OpenStudy (anonymous):

Help me find the volume of the following cones: 1. Height 5 cm, area of base 15m² 2. Height 14 cm, radius of base 6 cm 3. Height 24 cm, diameter of base 7 cm 4. Height 28 cm, circumference of base 132 cm 5. Height 12 cm, area of base 3950cm²

OpenStudy (anonymous):

use volume=1/3(pi r^2 h)

OpenStudy (anonymous):

that is the correct formula so area in #1 is pi r^2h for #3 r= (1/2)d for 4 c=2 pi r A = pi r^2

OpenStudy (amistre64):

thats alotta cones ....

OpenStudy (anonymous):

maybe they are openening an ice cream parlor and are looking into the size of their family size ice cream cones :)